OpenStudy (anonymous):

a square scarf is folded in half to form a rectangle , if the resulting rectangle has a perimeter of 15 inches , what are the area and perimeter of the square scarf ?

OpenStudy (dominusscholae):

Let s be the side of the square scarf. When we fold the scarf in half, we form a rectangle of length l= s. With the other side formed, the side s is halved, so the width of the new rectangle is s/2. Since the perimeter of a rectangle is 2(l+w), the perimeter, in terms of s, is 2(s+s/2)=2(3/2)s=3s. Setting this equal to 15 yields s=5. But the perimeter of a square is 4s, so the answer is 4(5)=20.